A group consists of 5 males and 6 females. Four people are selected from the group.
1. In how many ways can we choose 4 people from those 11?

2. In how many ways can we choose 4 females from the 6 females?

3. What is the probability that we get all 4 females when we choose 4 from the whole group?

4. What is the probability that we get at least 1 male when we choose 4 from the whole group?

Answer by stanbon(57361)

A group consists of 5 males and 6 females. Four people are selected from the group.
1. In how many ways can we choose 4 people from those 11?
11C4 = [11*10*9*8]/[1*2*3*4] = 330
------------------------------------------
2. In how many ways can we choose 4 females from the 6 females?
6C4 = 6C2 = [6*5]/[1*2] = 15
---------------------------------------
3. What is the probability that we get all 4 females when we choose 4 from the whole group?
15/330
-----------------
4. What is the probability that we get at least 1 male when we choose 4 from the whole group?
P(at least one male) = 1 - P(no male in choosing 4) = 1 - 15/330 = 315/330
=========================
Cheers,
Stan H.